During his training as a Canadian astronaut, David Saint-Jacques was submitted to accelerations in a horizontal "centrifuge". Such a device is considered to have a radius of 5.0 m and can generate a centripetal acceleration equivalent to 7 times the acceleration gravitational average on Earth.

At how many turn per minute should the rotation be made to have such an acceleration?
Answer #1

a = 7 g = 7*9.8 = 68.6 m/s^2

r w^2 = 68.6

5* w^2 = 68.6

w = 3.704 rad/s

w = 3.704* 60/(2* pi)

w = 35.39 turns/ min

==========

Comment in case any doubt, will reply for sure. Goodluck